# Lizhang Chen*†‡ Jonathan Li*† Kaizhao Liang Baiyu Su

### v1

@@ Appendix A Notation and Definitions
$\left\lVert\cdot\right\rVert_{p}$ denotes the $\ell_{p}$ norm for $p\in[1,\infty]$. $\left\langle\cdot,\cdot\right\rangle$ denotes the standard inner product on $\mathbb{R}^{d}$. $\mathbf{x}]_{i}$ denotes the $i^{\text{th}}$ entry of a vector $\mathbf{x}$. $\textbf{{diag}}\left(\mathbf{x}\right)$ denotes the diagonal matrix with diagonal entries given by $\mathbf{x}$. $\mathbb{I}(\mathbf{x}\geq\mathbf{0})$ denotes the indicator tensor that is $1$ in a coordinate if $\mathbf{x}$ is nonnegative in that coordinate and 0 otherwise. If $\mathcal{K}:\mathbb{R}^{d}\to\mathbb{R}$ is convex, we let $\partial\mathcal{K}(\mathbf{x})$ denote the set of subgradients of $\mathcal{K}$ at $\mathbf{x}$ and overload $\nabla\mathcal{K}(\mathbf{x})$ to denote an element of $\partial\mathcal{K}(\mathbf{x})$.
@@ D.4 Adam
Appendix E Deferred Proofs
We assume the setting of Theorem 1.
@@ Appendix E Deferred Proofs
It suffices to work in an arbitrary coordinate $i$. Let $m:=[\widehat{\mathbf{m}}_{t}]_{i}$, $v:=[\widehat{\mathbf{v}}_{t}]_{i}$, and $g_{t}:=[\mathbf{g}_{t}]_{i}$. By expanding the update rule for $m$ and $v$, we obtain
@@
Lemma 2.
For all $t\in\mathbb{N}$, $\left\lVert\mathbf{x}_{t}\right\rVert_{\infty}\leq R$, $\left\lVert\widehat{\mathbf{m}}_{t}\right\rVert_{\infty}\leq G$, and $\left\lVert\widehat{\mathbf{v}}_{t}\right\rVert_{\infty}\leq G^{2}$ for some constants $R$ and $G$. We can choose $G\geq 1$ without loss of generality.
Proof.
Since the iterates are bounded and $f$ is $L$-smooth by Assumption 1, there exist constants $R$ and $G$ such that $\left\lVert\mathbf{x}_{t}\right\rVert_{\infty}\leq R$ and $\left\lVert\nabla f(\mathbf{x}_{t})\right\rVert_{\infty}\leq G$ for all $t\in\mathbb{N}$. It follows that $\left\lVert\widehat{\mathbf{m}}_{t}\right\rVert_{\infty}\leq G$ and $\left\lVert\widehat{\mathbf{v}}_{t}\right\rVert_{\infty}\leq G^{2}$. ∎
@@
By smoothness, Lemma 1, and Lemma 2, we have
@@
$\mathbb{E}[-\left\langle\nabla f(\mathbf{x}_{t}),\mathbb{I}(\mathbf{m}_{t}\mathbf{x}_{t}\geq\mathbf{0})\mathbf{x}_{t}\right\rangle]\leq-\left\lVert(\nabla f(\mathbf{x}_{t})\mathbf{x}_{t})^{+}\right\rVert_{1}+\beta_{1}^{t}GRd+\frac{\beta_{1}\eta LRd(C+\lambda R)}{1-\beta_{1}}+\frac{\sigma Rd}{\sqrt{n_{\mathrm{batch}}(1+\beta_{1})}}.$
@@
$\mathbb{E}[-\left\langle\nabla f(\mathbf{x}_{t}),\mathbb{I}(\mathbf{m}_{t}\mathbf{x}_{t}\geq\mathbf{0})\mathbf{x}_{t}\right\rangle]\leq-\left\lVert(\nabla f(\mathbf{x}_{t})\mathbf{x}_{t})^{+}\right\rVert_{1}+\beta_{1}^{t}GRd+\frac{\beta_{1}\eta LRd(C+\lambda R)}{1-\beta_{1}}+\frac{\sigma Rd}{\sqrt{n_{\mathrm{batch}}(1+\beta_{1})}},$
@@
See 1
Proof.
@@
Taking the expectation of (15) and applying Lemmas 1, 2, 4, and 6,
@@
$\quad+\eta\lambda\left(-\left\lVert(\nabla f(\mathbf{x}_{t})\mathbf{x}_{t})^{+}\right\rVert_{1}+\beta_{1}^{t}GRd+\frac{\beta_{1}\eta LRd(C+\lambda R)}{1-\beta_{1}}+\frac{\sigma Rd}{\sqrt{n_{\mathrm{batch}}(1+\beta_{1})}}\right)$
$\quad+\frac{\eta^{2}L}{2}(C^{2}d+\lambda^{2}R^{2}d).$
Rearranging (16), using $1-\beta_{1}^{t}\leq 1$ and $G\geq 1$, summing over $T$ iterations, and dividing both sides by $T$ gives
@@
$\quad+\frac{\beta_{1}\eta\lambda LRd(C+\lambda R)(G+\epsilon)}{1-\beta_{1}}+\frac{\eta L(G+\epsilon)}{2}(C^{2}d+\lambda^{2}R^{2}d)$
@@
$K_{3}$ $:=\frac{\beta_{1}Ld(C+\lambda R)(G+\epsilon)}{1-\beta_{1}}\left(\frac{G}{\epsilon}+\lambda R\right)+\frac{1}{2}Ld(C^{2}+\lambda^{2}R^{2})(G+\epsilon)$
@@ Appendix G Additional Experiment Results
This section provides supplementary experimental analyses that further characterize the behavior of cautious weight decay (CWD) across different optimizers and training dynamics. We present detailed visualizations of the mask activation patterns (Figure 6), showing how the fraction of parameters receiving weight decay evolves during training for both AdamW and Muon optimizers. Additionally, we include comprehensive loss and accuracy curves for all three optimizers (AdamW, Lion, and Muon) across model scales from  [^1] M to 2.3B parameters (Figures 810), demonstrating consistent improvements with CWD. Finally, Figure 12 tracks the evolution of parameter norms throughout training, revealing that CWD maintains stable regularization comparable to standard weight decay while achieving superior performance. These results collectively illustrate that CWD’s selective application of weight decay leads to more effective optimization without compromising training stability.
@@
Refer to captionRefer to captionRefer to caption
Figure 7: Training dynamics for the 986M-parameter Gemma model.

### v2

@@ 6 Conclusion
Acknowledgments
This work was supported in part by the Institute for Foundations of Machine Learning (IFML) and the Office of Naval Research (ONR) under Grant No. N00014-25-1-2354.
@@ Appendix A Notation and Definitions
$\left\lVert\cdot\right\rVert_{p}$ denotes the $\ell_{p}$ norm for $p\in[1,\infty]$. $\left\langle\cdot,\cdot\right\rangle$ denotes the standard inner product on $\mathbb{R}^{d}$. $\mathbf{x}]_{i}$ denotes the $i^{\text{th}}$ entry of a vector $\mathbf{x}$. $\textbf{{diag}}\left(\mathbf{x}\right)$ denotes the diagonal matrix with diagonal entries given by $\mathbf{x}$. $\mathbb{I}(\mathbf{x}\geq\mathbf{0})$ denotes the indicator tensor that is 11 in a coordinate if $\mathbf{x}$ is nonnegative in that coordinate and 0 otherwise. If $\mathcal{K}:\mathbb{R}^{d}\to\mathbb{R}$ is convex, we let $\partial\mathcal{K}(\mathbf{x})$ denote the set of subgradients of $\mathcal{K}$ at $\mathbf{x}$ and overload $\nabla\mathcal{K}(\mathbf{x})$ to denote an element of $\partial\mathcal{K}(\mathbf{x})$.
@@ B.4 Muon
B.5 Adam
Algorithm 6 Adam with cautious weight decay
3:  repeat
7:   $\widehat{\mathbf{m}}_{t}\leftarrow(1-\beta_{1}^{t})^{-1}\mathbf{m}_{t}$
8:   $\widehat{\mathbf{v}}_{t}\leftarrow(1-\beta_{2}^{t})^{-1}\mathbf{v}_{t}$
[…]
10:   $t\leftarrow t+1$
11:  until stopping criterion is met
12:  return optimized parameters $\mathbf{x}_{t}$
@@ D.4 Adam
Appendix E Convergence Rate of Adam with Cautious Weight Decay
In this section, we show that under the following assumptions, Adam with cautious weight decay (Algorithm 6) achieves a convergence rate on the squared gradient norm and an additional stationarity measure.
Assumption 1 (Smoothness).
$f$ is $L$-smooth and lower bounded by a finite constant $f^{\star}$.
Assumption 2 (Bounded variance).
The stochastic gradient $\mathbf{g}_{t}$ satisfies
$\mathbb{E}[\mathbf{g}_{t}\mid\mathbf{x}_{t}]=\nabla f(\mathbf{x}_{t})\quad\text{and}\quad\textup{Var}(\mathbf{g}_{t})=\mathbb{E}\left[\left\lVert\mathbf{g}_{t}-\nabla f(\mathbf{x}_{t})\right\rVert_{2}^{2}\mid\mathbf{x}_{t}\right]\leq\frac{\sigma^{2}}{n_{\mathrm{batch}}},$
where $\sigma$ is a constant and $n_{\mathrm{batch}}$ denotes the batch size.
Assumption 3 (Bounded iterates and bounded gradients).
There exist constants $R$ and $G$ such that $\left\lVert\mathbf{x}_{t}\right\rVert_{\infty}\leq R$ and $\left\lVert\mathbf{g}_{t}\right\rVert_{\infty}\leq G$ a.s. for all $t\in\mathbb{N}$.
Assumptions 1 and 2 are standard and often used in the analysis of stochastic gradient algorithms [22, 3, 15, 1]. Assumption 3 can be justified using the Lyapunov function (8) if $f$ is additionally assumed to be coercive, since a Robbins–Siegmund argument with sufficiently small $\eta$ shows that the optimizer states remain in a compact sublevel set of $\mathcal{H}$ a.s. For the sake of clarity, here we take it as an explicit assumption. Similar assumptions have often been used for the analysis of Adam-style algorithms [37, 61, 83, 12, 15, 8].
Theorem 1.
Under Assumptions 1, 2, and 3, let $0\leq\beta_{1}\leq\beta_{2}<1$, $\lambda\geq 0$, $\epsilon>0$, and $\eta_{t}=\eta>0$, and suppose $\mathbf{x}_{t}$ is updated using Algorithm 6. Then for all $T\in\mathbb{N}$,
where $K_{1}$, $K_{2}$, $K_{3}$, and $K_{4}$ are constants depending only on $L$, $R$, $G$, $d$, $\epsilon$, $\lambda$, $\beta_{1}$, $\beta_{2}$, and $f(\mathbf{x}_{1})-f^{\star}$.
Remark 3.
Remark 4.
In the setting of Theorem 1, let $T\in\mathbb{N}$ and $\eta=\Theta\left(\frac{1}{\sqrt{T}}\right)$. Then
$\frac{1}{T}\sum_{t\in[T]}\mathbb{E}\left[\left\lVert\nabla f(\mathbf{x}_{t})\right\rVert_{2}^{2}+\lambda\left\lVert(\nabla f(\mathbf{x}_{t})\mathbf{x}_{t})^{+}\right\rVert_{1}\right]=O\left(\frac{1}{\sqrt{T}}+\frac{\sigma}{\sqrt{n_{\mathrm{batch}}}}\right).$
An $O(T^{-\frac{1}{2}})$ bound can be obtained by making the unrealistic assumption $n_{\mathrm{batch}}=\Theta(T)$. However, even without this assumption, the stated bound is of theoretical interest. For additional discussion, see [5, 83, 10].
@@ Appendix E Convergence Rate of Adam with Cautious Weight Decay
It suffices to work in an arbitrary coordinate $i$. Let $m:=[\widehat{\mathbf{m}}_{t}]_{i}$, $v:=[\widehat{\mathbf{v}}_{t}]_{i}$, and $g_{t}:=[\mathbf{g}_{t}]_{i}$. By expanding the update rules for $m$ and $v$, we obtain
@@
By smoothness, Lemma 1, and Assumption 3, we have
@@
$\mathbb{E}[-\left\langle\nabla f(\mathbf{x}_{t}),\mathbb{I}(\mathbf{m}_{t}\mathbf{x}_{t}\geq\mathbf{0})\mathbf{x}_{t}\right\rangle]\leq-\mathbb{E}\left[\left\lVert(\nabla f(\mathbf{x}_{t})\mathbf{x}_{t})^{+}\right\rVert_{1}\right]+\beta_{1}^{t}GRd+\frac{\beta_{1}\eta LRd(C+\lambda R)}{1-\beta_{1}}+\frac{\sigma Rd}{\sqrt{n_{\mathrm{batch}}(1+\beta_{1})}}.$
@@
$\mathbb{E}[-\left\langle\nabla f(\mathbf{x}_{t}),\mathbb{I}(\mathbf{m}_{t}\mathbf{x}_{t}\geq\mathbf{0})\mathbf{x}_{t}\right\rangle]\leq-\mathbb{E}\left[\left\lVert(\nabla f(\mathbf{x}_{t})\mathbf{x}_{t})^{+}\right\rVert_{1}\right]+\beta_{1}^{t}GRd+\frac{\beta_{1}\eta LRd(C+\lambda R)}{1-\beta_{1}}+\frac{\sigma Rd}{\sqrt{n_{\mathrm{batch}}(1+\beta_{1})}},$
@@
We are now ready to prove Theorem 1.
Proof of Theorem 1.
@@
Taking the expectation of (15) and applying Lemmas 1, 3, and 5,
@@
$\quad+\eta\lambda\left(-\mathbb{E}\left[\left\lVert(\nabla f(\mathbf{x}_{t})\mathbf{x}_{t})^{+}\right\rVert_{1}\right]+\beta_{1}^{t}GRd+\frac{\beta_{1}\eta LRd(C+\lambda R)}{1-\beta_{1}}+\frac{\sigma Rd}{\sqrt{n_{\mathrm{batch}}(1+\beta_{1})}}\right)$
$\quad+\eta^{2}L(C^{2}d+\lambda^{2}R^{2}d).$
We can assume $G\geq\frac{1}{1-\beta_{1}}$ without loss of generality. Rearranging (16), using $1-\beta_{1}^{t}\leq 1$ and $(1-\beta_{1}^{t})(G+\epsilon)\geq 1$, summing over $T$ iterations, and dividing both sides by $T$ gives
@@
$\quad+\frac{\beta_{1}\eta\lambda LRd(C+\lambda R)(G+\epsilon)}{1-\beta_{1}}+\eta L(G+\epsilon)(C^{2}d+\lambda^{2}R^{2}d)$
@@
$K_{3}$ $:=\frac{\beta_{1}Ld(C+\lambda R)(G+\epsilon)}{1-\beta_{1}}\left(\frac{G}{\epsilon}+\lambda R\right)+Ld(C^{2}+\lambda^{2}R^{2})(G+\epsilon)$
@@ Appendix G Additional Experiment Results
This section provides supplementary experimental analyses that further characterize the behavior of cautious weight decay (CWD) across different optimizers and training dynamics. We present detailed visualizations of the mask activation patterns (Figure 7), showing how the fraction of parameters receiving weight decay evolves during training for both AdamW and Muon optimizers. Additionally, we include comprehensive loss and accuracy curves for all three optimizers (AdamW, Lion, and Muon) across model scales from  [^1] M to 2.3B parameters (Figures 810), demonstrating consistent improvements with CWD. Finally, Figure 13 tracks the evolution of parameter norms throughout training, revealing that CWD maintains stable regularization comparable to standard weight decay while achieving superior performance.
@@
Refer to captionRefer to captionRefer to caption
Figure 11: Training dynamics for the 986M-parameter Gemma model.